# Robust targeted exploration for systems with non-stochastic disturbances

### v2

@@ 4.1 Sufficient conditions for exploration
[venkatasubramanian2023towards, Theorem 4] Suppose $\Phi$ and $X$ satisfy
@@
In order to compute the amplitudes of the exploration inputs $U_{\mathrm{e}}$, a few challenges need to be addressed. Note that Inequality (7) depends on $X$ and $\Phi$ quadratically, which further depend on the amplitudes of the exploration inputs $U_{\mathrm{e}}$ (6), as well as the disturbance $w$. Furthermore, since the true dynamics $A_{\mathrm{tr}},\,B_{\mathrm{tr}}$ are uncertain, the linear mapping from the input sequence to the state sequence is not known.
@@ 4.2 Bounds based on the theory of spectral lines
In what follows, we address the aforementioned issues by determining sufficient conditions for targeted exploration in terms of the spectral content of $X$ and $\Phi$ based on the theory of spectral lines. Given $u_{k}$ as in (6), $x_{k}$ has $L$ spectral lines from 0 to T1T-1 at distinct frequencies $\omega_{i}\in\Omega_{T}$, $i=1,\dots,L$ with amplitudes [sarker2020parameter, Lemma 1]:
@@
The transient error in the amplitude of a spectral line $\bar{x}_{\mathrm{err}}(\omega_{i})$ decays uniformly (with rate $\frac{1}{\sqrt{T}}$) to 0 as $T\to\infty$ (cf. Assumption 4, [ljung1999system, Theorem 2.1]). To simplify the exposition, we will assume that the transient error can be neglected.
@@
$\bar{Z}=$ $\begin{bmatrix}D_{\mathrm{des}}^{\frac{1}{2}\top}\left((\bar{X}_{\mathrm{u}}+\bar{X}_{\mathrm{w}})^{\mathsf{H}}\otimes I_{n_{\phi}}\right)\\ \left((\bar{\Phi}_{\mathrm{u}}+\bar{\Phi}_{\mathrm{w}})\otimes I_{n_{\mathrm{x}}}\right)\otimes I_{n_{\phi}}\end{bmatrix}$
$=$ $\underbrace{\begin{bmatrix}D_{\mathrm{des}}^{\frac{1}{2}\top}\left(\left(\mathbf{1}_{L}^{\top}U_{\mathrm{e}}^{\top}V_{\mathrm{x,tr}}^{\mathsf{H}}\right)\otimes I_{n_{\phi}}\right)\\ \left(V_{\mathrm{\phi,tr}}U_{\mathrm{e}}\right)\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}}_{=:\bar{Z}_{\mathrm{u}}}$
$+\underbrace{\begin{bmatrix}D_{\mathrm{des}}^{\frac{1}{2}\top}\left(\bar{X}_{\mathrm{w}}^{\mathsf{H}}\otimes I_{n_{\phi}}\right)\\ \bar{\Phi}_{\mathrm{w}}\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}}_{=:\bar{Z}_{\mathrm{w}}}.$ (33)
@@ 4.3 Bounds on transfer matrices
using $\theta_{\mathrm{tr}}\in\mathbf{\Theta}_{0}$ (cf. Assumption 2). Conditions (38) and (39) are LMIs, and hence matrices $\tilde{\Gamma}_{\mathrm{x}},\tilde{\Gamma}_{\phi}\succ 0$ and $\Gamma_{\phi},\Gamma_{\mathrm{\mathrm{x}}}\succ 0$ may be computed using robust control methods as shown in [venkatasubramanian2023sequential, Appendices A-B], or scenario methods as shown in [venkatasubramanian2023sequential, Appendix C]. Utilizing (39), we derive bounds on $\bar{\Phi}_{\mathrm{w}}$, $\bar{X}_{\mathrm{w}}$ and $\bar{Z}_{\mathrm{w}}$ in the following lemma.
@@
$\bar{Z}_{\mathrm{w}}\bar{Z}_{\mathrm{w}}^{\mathsf{H}}$ $\preceq\bar{W}_{Z}:=\left(\frac{\gamma_{\mathrm{w}}}{T}\|\Gamma_{\mathrm{x}}\|\|D_{\mathrm{des}}\|+\frac{\gamma_{\mathrm{w}}}{T}\|\Gamma_{\phi}\|\right)I_{(n_{\phi}+n_{\mathrm{x}}n_{\phi}^{2})}.$ (41)
The proof of Lemma 10 is provided in Appendix B. In what follows, given the bound on transfer matrices, we derive spectral lines-based sufficient conditions for targeted exploration.
@@ 4.4 Sufficient conditions for targeted exploration based on the theory of spectral lines
{pf}
Starting from Inequality (11), we have
[…]
Note that $\bar{\Phi}_{\mathrm{u}}$ and $\bar{Z}_{\mathrm{u}}$ depend linearly on the decision variable $U_{\mathrm{e}}$. Determining a lower bound based on the Inequality (11) results in non-convex constraints in $U_{\mathrm{e}}$. To overcome this problem, we utilize a convex relaxation procedure.
@@ 4.5 Convex relaxation
The following lemma is utilized to make Inequality (11) linear in the decision variable $U_{\mathrm{e}}$.
Lemma 12.
For any matrices $M\in\mathbb{C}^{n\times m}$ and $N\in\mathbb{C}^{n\times m}$, we have
$MM^{\mathsf{H}}\succeq MN^{\mathsf{H}}+NM^{\mathsf{H}}-NN^{\mathsf{H}}.$ (44)
{pf}
[…]
@@
{pf}
From Lemma 12, we have
[…]
@@ 4.6 Exploration SDP
The proof of Theorem 14 is provided in Appendix C. The key idea of the proof is the application of robust control tools, in particular, the matrix S-lemma [boyd2004convex, vanwaarde2022noisy], to account for parametric uncertainty. In particular, feasibility of LMIs (49b) and (49c) requires $(\tilde{\Gamma}_{\phi}-\hat{V}_{\mathrm{\phi}}\hat{V}_{\mathrm{\phi}}^{\mathsf{H}})\prec 0$, which holds if the initial uncertainty is sufficiently small. Consequently, we can pose the exploration problem of designing exploration inputs that excite the system with minimal energy to obtain an estimate $\hat{\theta}_{T}$ that satisfies (4) using the following SDP:
@@
$\hat{U}$ $=U_{\mathrm{e}}^{*},$
$\hat{Z}$ $=\begin{bmatrix}D_{\mathrm{des}}^{\frac{1}{2}\top}\left(\left(\mathbf{1}_{L}^{\top}U_{\mathrm{e}}^{*\top}\hat{V}_{\mathrm{x}}^{\mathsf{H}}\right)\otimes I_{n_{\phi}}\right)\\ \left(\hat{V}_{\mathrm{\phi}}U_{\mathrm{e}}^{*}\right)\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}$ (52)
@@
(Efficient implementation and conservatism) If $D_{\mathrm{des}}$ is a scaled identity matrix, i.e., $D_{\mathrm{des}}=cI_{n_{\phi}}$ with $c>0$, the Kronecker product with $I_{n_{\phi}}$ in LMIs (49a) and (49b) can be factored out. Hence, the dimensions of LMIs (49a) and (49b) can be reduced by a factor of $n_{\phi}$, i.e., to $(Ln_{\mathrm{u}}+1+n_{\mathrm{x}}n_{\phi})$, and to $(Ln_{\mathrm{u}}n_{\mathrm{x}}+1+n_{\mathrm{x}}n_{\phi})$, respectively. The resulting LMIs are equivalent to the LMIs (49a) and (49b) before factoring. Furthermore, LMIs (49a)-(49c) are derived by invoking the matrix S-lemma [van2020noisy]. Enforcing (48a)-(48c) through LMIs (49a)-(49c) does not introduce any conservatism because the satisfaction of LMIs (49a)-(49c) is equivalent to the satisfaction of (48a)-(48c) under conditions (36) and (38).
The overall targeted exploration strategy is summarized in Algorithm 1. The resulting strategy optimally excites the system with exploratory inputs (6) in order to determine model parameters up to a user defined closeness $D_{\mathrm{des}}$.
@@
3:Compute bounds $\tilde{\Gamma}_{\phi}$, $\tilde{\Gamma}_{\mathrm{x}}$ (38), and $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$ (39) via methods described in [venkatasubramanian2023sequential, Appendices A-C].
@@
5:Select initial candidates $\hat{Z}$ and $\hat{U}$ (4.6).
@@ 4.7 Discussion
Summary - proposed approach: The proposed targeted exploration strategy outlined in Algorithm 1 yields a multi-sine exploration input with minimal input energy to generate data from which estimates of uncertain parameters can be derived with a desired error bound (4). The frequencies $\omega_{i}$ of the multi-sine input (6) are predetermined, enabling intuitive tuning based on prior knowledge about the system. The proposed approach assumes energy-bounded disturbances, as commonly considered for data-driven models [fogel1979system, van2020noisy, berberich2022combining]. As the main result, the data-dependent uncertainty bound in Lemma 6 is utilized to derive sufficient conditions in the spectral content of the exploration inputs. The proposed exploration is targeted, as the optimized amplitudes at different frequencies impact both the magnitude and the shape/orientation of the remaining uncertainty after exploration. Furthermore, the proposed exploration strategy is robust, i.e., parametric uncertainty is accounted for by using robust control tools.
@@
Related works: The derived targeted exploration strategy is similar to, and inspired by [venkatasubramanian2023sequential, umenberger2019robust, ferizbegovic2019learning, barenthin2008identification], however, with a few crucial differences. A targeted exploration method is proposed in [barenthin2008identification] to identify parameters up to a desired accuracy. However, the conditions are not robust to uncertainty and hence iterative experiments are required in practice. In [ferizbegovic2019learning], the proposed targeted exploration strategy assumes independent data, and hence lacks applicability to correlated time-series data from a single trajectory. Furthermore, the methods in [ferizbegovic2019learning] and [umenberger2019robust] do not yield any guarantees for exploration since the uncertainty bounds are approximated in a heuristic way. The method in [venkatasubramanian2023sequential] robustly accounts for parametric uncertainties and provides an a priori guaranteed bound on the uncertainty after exploration. However, all the methods [barenthin2008identification, ferizbegovic2019learning, umenberger2019robust, venkatasubramanian2023sequential] assume i.i.d. Gaussian disturbances. In contrast to these methods, we consider energy-bounded disturbances without assumptions on the distribution or independence of the disturbances. This encompasses a broader class of uncertainties, including those arising from unmodeled dynamics or nonlinearities, and enables the development of a targeted exploration strategy with guarantees. In our proposed targeted exploration method, we quantify and guarantee a priori uncertainty bounds on the parameters obtained from an experiment. Similar to [venkatasubramanian2023sequential], we provide a priori guarantees by robustly accounting for the impact of the uncertain model parameters and disturbances. A notable difference from [venkatasubramanian2023sequential] is the appearance of the term $ZZ^{\top}$ in Theorem 7, which cannot be directly addressed with standard robust control tools and presented significant additional challenges in the present paper.
Application to dual control: The proposed targeted exploration strategy may be utilized to design a robust dual control strategy which guarantees a desired performance for the closed loop after exploration [barenthin2008identification, venkatasubramanian2023sequential]. This can be achieved by co-designing the targeted exploration problem (4.6) with a robust gain-scheduled controller in order to account for the changes in uncertainty during the process of exploration [venkatasubramanian2023sequential]. Such a joint design of a dual controller also allows optimizing closed-loop performance after exploration, ensuring the exploration process reduces only the necessary uncertainty to achieve the desired performance within exploration energy constraints.
@@ 5 Numerical Example
In this section, we demonstrate the applicability of the proposed targeted exploration strategy to a nonlinear system using a numerical example. Numerical simulations [^1] The source code for the simulations is available at https:////github.com//jananivenkatasubramanian//NonstochTE were performed on MATLAB using CVX [cvx] in conjunction with the solver MOSEK.
@@
and
$F_{i}^{\mathrm{nl}}$ $=\beta_{i}\tanh(\alpha_{i}\dot{p}_{i}),$ (54)
with positions $p_{i}$, velocities $\dot{p}_{i}$, masses $m_{i}>0$, spring constants $k_{i}\geq 0$, damping coefficients $d_{i}\geq 0$, nonlinear Coulomb frictions $F_{i}^{\mathrm{nl}}$ with constants $\alpha_{i},\,\beta_{i}\geq 0$, $i\in\{1,2\}$, and input $u\in\mathbb{R}$. We consider $x=[p_{1},\dot{p}_{1},p_{2},\dot{p}_{2}]^{\top}\in\mathbb{R}^{4}$ and an Euler discretization with sampling period $T_{s}=0.5$. The resulting system dynamics correspond to our setup (2) with the disturbances $w_{k}=\left[0,\frac{F_{1}^{\mathrm{nl}}(x_{2})}{m_{1}},0,\frac{F_{1}^{\mathrm{nl}}(x_{4})}{m_{2}}\right]^{\top}$. The disturbances satisfy Assumption 1 with $\gamma_{\mathrm{w}}=T\left(\frac{\beta_{1}^{2}}{m_{1}^{2}}+\frac{\beta_{2}^{2}}{m_{2}^{2}}\right)$. In this example, we consider the following true values for the masses, spring constants and damping coefficients: $m_{1}=1$, $m_{2}=2$, $k_{1}=1$, $k_{2}=1.5$, $d_{1}=0.5$ and $d_{2}=1.1$ (5). We select the coefficients of Coulomb friction as $\alpha_{1}=1$ and $\alpha_{2}=1$ in (54). Later, we vary $\beta_{1}$ and $\beta_{2}$ to study the effect of the disturbance bound $\gamma_{\mathrm{w}}$.
The goal of the proposed targeted exploration strategy is to achieve a desired error bound $D_{\mathrm{des}}^{-1}=10^{0}I_{n_{\phi}}$ on the parameters (4). Furthermore, we select $L=20$ equally-spaced frequencies $\omega_{i}\in\{0,0.05,0.1,0.15,...,0.95\}$. We set $T=100$ and $\epsilon=0.5$. In what follows, we analyse the effectiveness and conservatism of the proposed targeted exploration strategy.
Required input energy $\gamma_{\mathrm{e}}^{2}$ for different energy bounds $\gamma_{\mathrm{w}}$: To study the effect of the energy bound $\gamma_{\mathrm{w}}$ on the required input energy $\gamma_{\mathrm{e}}^{2}$ that ensures the exploration goal, we select the initial uncertainty level as $D_{0}^{-1}=10^{-4}I_{n_{\phi}}$ and the initial estimate as $\hat{\theta}_{0}=\theta_{\mathrm{tr}}+\frac{\|D_{0}\|^{\frac{1}{2}}}{\|\theta_{\mathrm{tr}}\|}\theta_{\mathrm{tr}}$, i.e., on the boundary of $\mathbf{\Theta}_{0}$ (cf. Assumption 2). We run five trials for the following energy bounds $\gamma_{\mathrm{w}}\in\{10^{-4},10^{-3},10^{-2},10^{-1},10^{0}\}$. Each trial comprises: $\mathrm{(i)}$ computing the corresponding constants $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$ using the scenario approach [venkatasubramanian2023sequential, Appendix C] with confidence level $\beta=10^{-10}$, $\mathrm{(ii)}$ executing Algorithm 1 to obtain the exploration inputs (6) and the required input energy $\gamma_{\mathrm{e}}^{2}$.
@@
From Figure 1, it can be observed that the input energy $\gamma_{\mathrm{e}}^{2}$ scales roughly linearly with the disturbance energy bound $\gamma_{\mathrm{w}}$. Furthermore, the input energy needs to be significantly larger than the disturbance energy to achieve the desired accuracy of the estimated parameters. In general, $\gamma_{\mathrm{e}}\to 0$ as $\gamma_{\mathrm{w}}\to 0$, i.e., the exploration input energy $\gamma_{\mathrm{e}}$ reduces to zero as the energy of the disturbances $\gamma_{\mathrm{w}}$ reduce to zero.
Conservatism related to the initial uncertainty bound $D_{0}$: To study the effect of the initial uncertainty bound $D_{0}^{-1}$, we select eight uncertainty levels $D_{0}\in\{10^{i}I_{n_{\phi}},5\times 10^{i}I_{n_{\phi}}\}$, i=4,,7i=4,...,7 and $\gamma_{\mathrm{w}}=1$. Based on the energy bound $\gamma_{\mathrm{w}}$, we compute $\beta_{1}=\beta_{2}=\frac{m_{1}m_{2}\sqrt{\gamma_{\mathrm{w}}}}{\sqrt{T(m_{1}^{2}+m_{2}^{2})}}$. Each trial comprises: $\mathrm{(i)}$ computing the initial estimate as $\hat{\theta}_{0}=\theta_{\mathrm{tr}}+\frac{\|D_{0}\|^{\frac{1}{2}}}{\|\theta_{\mathrm{tr}}\|}\theta_{\mathrm{tr}}$, $\mathrm{(ii)}$ computing the corresponding constants $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$ using the scenario approach [venkatasubramanian2023sequential, Appendix C] with confidence level $\beta=10^{-10}$, $\mathrm{(iii)}$ executing Algorithm 1 to obtain the exploration inputs (6), and generating a dataset by applying the exploration input. From the datasets, we compute the estimate of the parameters $\hat{\theta}_{T}$, and $\|G\cdot P\|$, the error bound guaranteed by Lemma 6 with the data-dependent covariance matrix $P$ (12) and scaling $G$ (14).
Figure 2: Illustration of the a posteriori guaranteed bound on the squared error of the parameters $\|G\cdot P\|$, in comparison with the desired bound on the squared error $\|D_{\mathrm{des}}^{-1}\|$ for different initial uncertainty bounds $D_{0}^{-1}$.
From Figure 2, it can be observed that the targeted exploration inputs achieve the exploration goal for all tested initial uncertainty levels $D_{0}$ and energy bound $\gamma_{\mathrm{w}}$, i.e., the guaranteed error bound on the parameters $\|G\cdot P\|$ is lower than the desired error bound $\|D_{\mathrm{des}}^{-1}\|~=~10^{0}$. The guaranteed error bound is significantly lower than the desired error bound due to the inherent conservatism of the proposed strategy which is derived by utilizing worst case bounds on the transfer matrices $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$, and worst case bounds $\bar{W}_{\phi}$, $\bar{W}_{Z}$. This conservatism decreases as the initial uncertainty is reduced, or equivalently, as $D_{0}$ increases. This occurs because smaller initial uncertainty results in tighter and more accurate bounds on $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$, allowing for more efficient and targeted exploration.

### v3

@@ 4.1 Sufficient conditions for exploration
[28, Theorem 4] Suppose $\Phi$ and $X$ satisfy
@@ 4.2 Bounds based on the theory of spectral lines
In what follows, we address the aforementioned issues by determining sufficient conditions for targeted exploration in terms of the spectral content of $X$ and $\Phi$ based on the theory of spectral lines. Given $u_{k}$ as in (6), $x_{k}$ has $L$ spectral lines from 0 to T1T-1 at distinct frequencies $\omega_{i}\in\Omega_{T}$, $i=1,\dots,L$ with amplitudes [21, Lemma 1]:
@@
The transient error in the amplitude of a spectral line $\bar{x}_{\mathrm{err}}(\omega_{i})$ decays uniformly (with rate $\frac{1}{\sqrt{T}}$) to 0 as $T\to\infty$ (cf. Assumption 4, [15, Theorem 2.1]). To simplify the exposition, we will assume that it can be neglected [21].
@@
$\bar{Z}=$ $\begin{bmatrix}D_{\mathrm{des}}^{\frac{1}{2}\top}((\bar{X}_{\mathrm{u}}+\bar{X}_{\mathrm{w}})^{\mathsf{H}}\otimes I_{n_{\phi}})\\ ((\bar{\Phi}_{\mathrm{u}}+\bar{\Phi}_{\mathrm{w}})\otimes I_{n_{\mathrm{x}}})\otimes I_{n_{\phi}}\end{bmatrix}$
$=$ $\underbrace{\begin{bmatrix}D_{\mathrm{des}}^{\frac{1}{2}\top}((\mathbf{1}_{L}^{\top}U_{\mathrm{e}}^{\top}V_{\mathrm{x,tr}}^{\mathsf{H}})\otimes I_{n_{\phi}})\\ (V_{\mathrm{\phi,tr}}U_{\mathrm{e}})\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}}_{=:\bar{Z}_{\mathrm{u}}}+\underbrace{\begin{bmatrix}D_{\mathrm{des}}^{\frac{1}{2}\top}(\bar{X}_{\mathrm{w}}^{\mathsf{H}}\otimes I_{n_{\phi}})\\ \bar{\Phi}_{\mathrm{w}}\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}}_{=:\bar{Z}_{\mathrm{w}}}.$ (30)
@@ 4.3 Bounds on transfer matrices
using $\theta_{\mathrm{tr}}\in\mathbf{\Theta}_{0}$ (cf. Assumption 2). Conditions (35) and (36) are LMIs, and hence matrices $\tilde{\Gamma}_{\mathrm{x}},\tilde{\Gamma}_{\phi}\succ 0$ and $\Gamma_{\phi},\Gamma_{\mathrm{\mathrm{x}}}\succ 0$ may be computed using robust control methods as shown in [26, Appendices A-B], or scenario methods as shown in [26, Appendix C]. Utilizing (36), we derive bounds on $\bar{\Phi}_{\mathrm{w}}$, $\bar{X}_{\mathrm{w}}$ and $\bar{Z}_{\mathrm{w}}$ in the following lemma.
@@
$\bar{Z}_{\mathrm{w}}\bar{Z}_{\mathrm{w}}^{\mathsf{H}}$ $\preceq\underbrace{\left(\frac{\gamma_{\mathrm{w}}}{T}\|\Gamma_{\mathrm{x}}\|\|D_{\mathrm{des}}\|+\frac{\gamma_{\mathrm{w}}}{T}\|\Gamma_{\phi}\|\right)I_{(n_{\phi}+n_{\mathrm{x}}n_{\phi}^{2})}}_{\bar{W}_{Z}:=}.$ (38)
The proof of Lemma 10 is provided in Appendix B. Given the bounds in Lemma 10, we next derive spectral lines-based sufficient conditions for targeted exploration.
@@ 4.4 Sufficient conditions for targeted exploration based on the theory of spectral lines
The proof of Proposition 11 is provided in Appendix C. Note that $\bar{\Phi}_{\mathrm{u}}$ and $\bar{Z}_{\mathrm{u}}$ depend linearly on the decision variable $U_{\mathrm{e}}$. Determining a lower bound based on the Inequality (11) results in non-convex constraints in $U_{\mathrm{e}}$. To overcome this problem, we utilize a convex relaxation procedure.
@@ 4.5 Convex relaxation
The proof of the proposition is provided in Appendix D. The bound derived in (12) is tight if $\hat{Z}=\bar{Z}_{\mathrm{u}}$. Since $\bar{Z}_{\mathrm{u}}$ comprises uncertain elements $V_{\mathrm{x,tr}}$, $V_{\mathrm{\phi,tr}}$ and the unknown decision variable $U_{\mathrm{e}}$, we consider a candidate $\hat{Z}$. Later, this relaxation is embedded an iterative process to reduce conservatism. In what follows, we utilize Inequality (12) to derive a condition linear in the decision variable $U_{\mathrm{e}}$ that ensures the exploration goal (4).
@@ 4.6 Exploration SDP
The proof of Theorem 13 is provided in Appendix E. The key idea of the proof is the application of robust control tools, in particular, the matrix S-lemma [5, 24], to account for parametric uncertainty. In particular, feasibility of LMIs (49b) and (49c) requires $(\tilde{\Gamma}_{\phi}-\hat{V}_{\mathrm{\phi}}\hat{V}_{\mathrm{\phi}}^{\mathsf{H}})\prec 0$, which holds if the initial uncertainty is sufficiently small. Consequently, we can pose the exploration problem of designing exploration inputs that excite the system with minimal energy to obtain an estimate $\hat{\theta}_{T}$ that satisfies (4) using the following SDP:
@@
$\hat{U}$ $=U_{\mathrm{e}}^{*},\;\hat{Z}$ $=\begin{bmatrix}D_{\mathrm{des}}^{\frac{1}{2}\top}\left(\left(\mathbf{1}_{L}^{\top}U_{\mathrm{e}}^{*\top}\hat{V}_{\mathrm{x}}^{\mathsf{H}}\right)\otimes I_{n_{\phi}}\right)\\ \left(\hat{V}_{\mathrm{\phi}}U_{\mathrm{e}}^{*}\right)\otimes I_{n_{\mathrm{x}}n_{\phi}}\end{bmatrix}$ (46)
@@
(Efficient implementation and conservatism) If $D_{\mathrm{des}}$ is a scaled identity matrix, i.e., $D_{\mathrm{des}}=cI_{n_{\phi}}$ with $c>0$, the Kronecker product with $I_{n_{\phi}}$ in LMIs (43a) and (43b) can be factored out. Hence, the dimensions of LMIs (43a) and (43b) can be reduced by a factor of $n_{\phi}$, i.e., to $(Ln_{\mathrm{u}}+1+n_{\mathrm{x}}n_{\phi})$, and to $(Ln_{\mathrm{u}}n_{\mathrm{x}}+1+n_{\mathrm{x}}n_{\phi})$, respectively. The resulting LMIs are equivalent to the LMIs (43a) and (43b) before factoring. Furthermore, LMIs (43a)-(43c) are derived by invoking the matrix S-lemma [24]. LMIs (43a)-(43c) are sufficient conditions for (42a)-(42c) under conditions (33) and (35). However, due to the imposed structural constraints due to Kronecker products, the LMIs (43a)-(43c) may, in general, be conservative.
Remark 15.
(Computational effort) The computational effort of Problem (4.6) scales polynomially with the state dimension $n_{\mathrm{x}}$, control input dimension $n_{\mathrm{u}}$, and the number of frequencies $L$. While this makes moderately-sized problems tractable, large-scale problems may become computationally demanding. In such cases, problem-specific structures and efficient solvers may be utilized to reduce computational demand [16].
The overall exploration strategy is summarized in Algorithm 1. The resulting strategy optimally excites the system with inputs (6) to determine model parameters up to a user defined closeness $D_{\mathrm{des}}$.
@@
3:Compute bounds $\tilde{\Gamma}_{\phi}$, $\tilde{\Gamma}_{\mathrm{x}}$ (35), and $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$ (36) via methods described in [26, Appendices A-C].
@@
5:Select initial candidates $\hat{Z}$ and $\hat{U}$ (46).
@@ 4.7 Discussion
Summary - proposed approach: The proposed targeted exploration strategy outlined in Algorithm 1 yields a multi-sine exploration input with minimal input energy to generate data from which estimates of uncertain parameters can be derived with a desired error bound (4). The frequencies $\omega_{i}$ of the multi-sine input (6) are predetermined, enabling intuitive tuning based on prior knowledge about the system. The proposed approach assumes energy-bounded disturbances, as commonly considered for data-driven models [9, 24, 2]. As the main result, the data-dependent uncertainty bound in Lemma 6 is utilized to derive sufficient conditions in the spectral content of the exploration inputs. The proposed exploration is targeted, as the optimized amplitudes at different frequencies impact both the magnitude and the shape/orientation of the remaining uncertainty after exploration. Furthermore, the proposed exploration strategy is robust, i.e., parametric uncertainty is accounted for by using robust control tools.
@@
Related works: The derived targeted exploration strategy is similar to, and inspired by [26, 23, 8, 1], however, with a few crucial differences. A targeted exploration method is proposed in [1] to identify parameters up to a desired accuracy. However, the conditions are not robust to uncertainty, requiring iterative experiments in practice. In [8], the proposed targeted exploration strategy assumes independent data, and hence lacks applicability to correlated time-series data from a single trajectory. Furthermore, the methods in [8] and [23] do not yield any guarantees for exploration since the uncertainty bounds are heuristically approximated. The method in [26] robustly accounts for parametric uncertainties and provides an a priori guaranteed bound on the uncertainty after exploration. However, all the methods [1, 8, 23, 26] assume i.i.d. Gaussian disturbances. In contrast to these methods, we consider energy-bounded disturbances without assumptions on the distribution or independence of the disturbances. This encompasses a broader class of uncertainties, including those arising from unmodeled dynamics or nonlinearities, and enables the development of a targeted exploration strategy with guarantees. In our proposed targeted exploration method, we quantify and guarantee a priori uncertainty bounds on the parameters obtained from a single experiment by robustly accounting for the impact of the uncertain model parameters and disturbances, similar to [26]. A notable difference to the theoretical analysis in [26] is the appearance of the term $ZZ^{\top}$ in Theorem 7, which cannot be directly addressed with standard robust control tools and presented significant additional challenges in this work.
Application to dual control: The proposed targeted exploration strategy can be integrated into a robust dual control framework to guarantee a desired closed-loop performance after exploration [1, 26]. This can be achieved by co-designing the targeted exploration problem (4.6) with a robust gain-scheduled controller that accounts for changes in uncertainty during exploration [26]. Such a joint design allows optimizing closed-loop performance after exploration, while ensuring exploration reduces only the necessary uncertainty to achieve the desired performance within exploration energy constraints.
@@ 5 Numerical Example
In this section, we demonstrate the applicability of the proposed targeted exploration strategy to a nonlinear system using a numerical example. Numerical simulations [^1] The source code for the simulations is available at https:////github.com//jananivenkatasubramanian//NonstochTE were performed on MATLAB using CVX [12] in conjunction with the solver MOSEK.
@@
$F_{i}^{\mathrm{nl}}$ $=\zeta_{i}\tanh(\alpha_{i}\dot{p}_{i}),$ (48)
with positions $p_{i}$, velocities $\dot{p}_{i}$, masses $m_{i}>0$, spring constants $k_{i}\geq 0$, damping coefficients $d_{i}\geq 0$, nonlinear Coulomb frictions $F_{i}^{\mathrm{nl}}$ with constants $\alpha_{i},\,\zeta_{i}\geq 0$, $i\in\{1,2\}$, and input $u\in\mathbb{R}$. We consider $x=[p_{1},\dot{p}_{1},p_{2},\dot{p}_{2}]^{\top}\in\mathbb{R}^{4}$ and an Euler discretization with sampling period $T_{s}=0.5$. The resulting system dynamics correspond to our setup (2) with the disturbances $w_{k}=\left[0,\frac{F_{1}^{\mathrm{nl}}(x_{2})}{m_{1}},0,\frac{F_{1}^{\mathrm{nl}}(x_{4})}{m_{2}}\right]^{\top}$. The disturbances satisfy Assumption 1 with $\gamma_{\mathrm{w}}=T\left(\frac{\zeta_{1}^{2}}{m_{1}^{2}}+\frac{\zeta_{2}^{2}}{m_{2}^{2}}\right)$. In this example, we consider the following true values for the masses, spring constants and damping coefficients: $m_{1}=1$, $m_{2}=2$, $k_{1}=1$, $k_{2}=1.5$, $d_{1}=0.5$ and $d_{2}=1.1$ (47). We select the coefficients of Coulomb friction as $\alpha_{1}=1$ and $\alpha_{2}=1$ in (48). Later, we vary $\zeta_{1}$ and $\zeta_{2}$ to study the effect of the disturbance bound $\gamma_{\mathrm{w}}$.
The goal of the proposed targeted exploration strategy is to achieve a desired error bound $D_{\mathrm{des}}^{-1}=10^{0}I_{n_{\phi}}$ on the parameters (4). Furthermore, we select $L=20$ equally-spaced frequencies $\omega_{i}\in\{0,0.05,0.1,0.15,...,0.95\}$. We set $T=100$ and $\epsilon=0.5$. Throughout the example, the constants $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$ are computed using the scenario approach [26, Appendix C] with confidence level $\beta=10^{-10}$. In what follows, we analyse the effectiveness and conservatism of the proposed targeted exploration strategy.
Required input energy $\gamma_{\mathrm{e}}^{2}$ for different energy bounds $\gamma_{\mathrm{w}}$: To study the effect of the energy bound $\gamma_{\mathrm{w}}$ on the required input energy $\gamma_{\mathrm{e}}^{2}$, we select the initial uncertainty level as $D_{0}^{-1}=10^{-4}I_{n_{\phi}}$ and the initial estimate as $\hat{\theta}_{0}=\theta_{\mathrm{tr}}+\frac{\|D_{0}\|^{\frac{1}{2}}}{\|\theta_{\mathrm{tr}}\|}\theta_{\mathrm{tr}}$, i.e., on the boundary of $\mathbf{\Theta}_{0}$ (cf. Assumption 2). We run five trials for the following energy bounds $\gamma_{\mathrm{w}}\in\{10^{-4},10^{-3},10^{-2},10^{-1},10^{0}\}$. Each trial comprises: $\mathrm{(i)}$ computing the corresponding constants $\Gamma_{\phi}$, $\Gamma_{\mathrm{x}}$, $\tilde{\Gamma}_{\phi}$ and $\tilde{\Gamma}_{\mathrm{x}}$, $\mathrm{(ii)}$ executing Algorithm 1 to obtain the exploration inputs (6) and the required input energy $\gamma_{\mathrm{e}}^{2}$.
@@
From Figure 1, it can be observed that the input energy $\gamma_{\mathrm{e}}^{2}$ scales roughly linearly with the disturbance energy bound $\gamma_{\mathrm{w}}$. In general, $\gamma_{\mathrm{e}}\to 0$ as $\gamma_{\mathrm{w}}\to 0$, i.e., the exploration input energy $\gamma_{\mathrm{e}}$ reduces to zero as the energy of the disturbances $\gamma_{\mathrm{w}}$ reduce to zero.